# Engineering Artificial Intelligence: Framework, Challenges, and Future Direction

### v1

@@ Abstract
Over the past ten years, the application of artificial intelligence (AI) and machine learning (ML) in engineering domains has gained significant popularity, showcasing their potential in data-driven contexts. However, the complexity and diversity of engineering problems often require the development of domain-specific AI approaches, which are frequently hindered by a lack of systematic methodologies, scalability, and robustness during the development process. To address this gap, this paper introduces the ”ABCDE” as the key elements of Engineering AI and proposes a unified, systematic engineering AI ecosystem framework, including eight essential layers, along with attributes, goals, and applications, to guide the development and deployment of AI solutions for specific engineering needs. Additionally, key challenges are examined, and nine future research directions are highlighted. By providing a comprehensive perspective, this paper aims to advance the strategic implementation of AI, fostering the development of next-generation engineering AI solutions.
@@ 2 Key Elements in Engineering AI: ABCDE
The foundational components of engineering AI can be categorized into the ”ABCDE” as shown in Figure 1, where the first three components—Algorithm, Big Data, and Computing Power—form the key elements of AI technology. (A. Algorithm) represents advanced AI and ML methods, (B. Big Data) involves collection, storage, and processing of engineering-specific data, and (C. Computing Power) refers to state-of-the-art computational infrastructure, including high-performance chips and cloud resources. Together, these elements build the foundation for AI-driven tasks in engineering. Beyond these ”ABC” components, the ”DE” pairs extend into three critical race tracks: (AI+) Infrastructure, ($\times$×AI) Platform, and (+AI) Product. 1) (AI+) Infrastructure track emphasizes physical and energy systems, with (D.1 Data Center) supporting AI computation and data storage, and (E.1 Electricity) powering AI infrastructure. 2) ($\times$×AI) Platform track focuses on industrial domain knowledge (D.2 Domain) and AI-enabled ecosystems (E.2 Ecosystem) development to create AI solutions for different engineering domains. 3) the (+AI) Product track bridges AI-powered devices (D.3 Device) with edge computing environments (E.3 Edge) to enable real-time processing and decision-making near the data source. Overall, the ABC components form the core technological foundation required to develop AI systems, while the DE components define where and how AI technologies enabled by ABC are deployed and utilized. The flow from (AI+) Infrastructure, to (×AI) Platform, and to (+AI) Product illustrates a progressive development path in engineering AI, where each track contributes a distinct role: enabling computation and energy support (Data Center, Electricity), integrating domain knowledge and ecosystems (Domain, Ecosystem), and deploying intelligent, embedded applications at the edge (Device, Edge). Collectively, these tracks define the practical landscape for engineering AI.
@@ 5 Future Direction
Figure 4 provides an overview of the emerging future directions in engineering AI. This perspective helps to better understand the key methodologies and advancements and identify critical areas for future research. The nine key research directions are outlined as follows:
Figure 4: Future directions in engineering AI research. Each subfigure is collected from its respective source: (1) Transfer Learning & Domain Adaptation [16]; (2) Stream-of-Quality & Stream-of-X [31]; (3) Similarity-Based Machine Learning [67]; (4) Digital Twin-Based Machine Learning [20]; (5) Large Language Model & Industrial Large Knowledge Model [33]; (6) Multi-Modal Machine Learning [18]; (7) Synthetic Data Generation [41]; (8) Foundation Model [10]; (9) Hybrid Physics-Based and Data-Driven Models [63].
@@
Synthetic data generation Moderate Moderate Academia
@@
Synthetic Data Generation: addresses the challenge of limited data availability or imbalanced datasets by creating artificial data [22, 41, 8]. In the future, researchers can focus on improving the realism of generated data and ensuring its alignment with real-world engineering conditions by leveraging generative adversarial networks (GANs), diffusion models, and cutting-edge generative AI techniques.

### v2

@@ Abstract
Over the past ten years, the application of artificial intelligence (AI) and machine learning (ML) in engineering domains has gained significant popularity, showcasing their potential in data-driven contexts. However, the complexity and diversity of engineering problems often require the development of domain-specific AI approaches, which are frequently hindered by a lack of systematic methodologies, scalability, and robustness during the development process. To address this gap, this paper introduces the ”ABCDE” as the key elements of Engineering AI and proposes a unified, systematic engineering AI ecosystem framework, including eight essential layers, along with attributes, goals, and applications, to guide the development and deployment of AI solutions for specific engineering needs. Additionally, key challenges are examined, and eight future research directions are highlighted. By providing a comprehensive perspective, this paper aims to advance the strategic implementation of AI, fostering the development of next-generation engineering AI solutions.
@@ 2 Key Elements in Engineering AI: ABCDE
The foundational components of engineering AI can be categorized into the ”ABCDE” as shown in Figure 1, where the first three components—Algorithm, Big Data, and Computing Power—form the key elements of AI technology. (A. Algorithm) represents advanced AI and ML methods, (B. Big Data) involves collection, storage, and processing of engineering-specific data, and (C. Computing) refers to state-of-the-art computational infrastructure, including high-performance chips and cloud resources. Together, these elements build the foundation for AI-driven tasks in engineering. Beyond these ”ABC” components, the ”DE” pairs extend into three key paths: (AI+) Infrastructure, ($\times$×AI) Platform, and (+AI) Product. 1) (AI+) Infrastructure path emphasizes physical and energy systems, with (D.1 Data Center) supporting AI computation and data storage, and (E.1 Electricity) powering AI infrastructure. 2) ($\times$×AI) Platform path focuses on industrial domain knowledge (D.2 Domain) and AI-enabled ecosystems (E.2 Ecosystem) development to create AI solutions for different engineering domains. 3) the (+AI) Product path bridges AI-powered devices (D.3 Device) with edge computing environments (E.3 Edge) to enable real-time processing and decision-making near the data source. While the ABC components form the core technological foundation required to develop AI systems, they influence each other dynamically. Advances in algorithms require greater computing power and more diverse data, while computational innovation and large amounts of domain-specific data drive the development of new algorithms. The DE components define where and how AI technologies enabled by ABC components are deployed and utilized. These paths define the practical landscape for engineering AI, where each path contributes a distinct role: enabling computation and energy support (Data Center, Electricity), integrating domain knowledge and ecosystems (Domain, Ecosystem), and deploying intelligent, embedded applications at the edge (Device, Edge).
@@ 4 Challenges of Engineering Artificial Intelligence
Figure 4: Future directions in engineering AI research. Each subfigure is collected from its respective source: (1) Transfer Learning & Domain Adaptation [34]; (2) Stream-of-Quality & Stream-of-X [28]; (3) Similarity-Based Machine Learning [32]; (4) Digital Twin-Based Machine Learning [26]; (5) Multi-Modal Machine Learning [38]; (6) Hybrid Physics-Based and Data-Driven Models [33]; (7) Large Language Model & Industrial Large Knowledge Model [30]; (8) Foundation Model [9].
@@ 5 Future Direction
Figure 4 provides an overview of the emerging future directions in engineering AI. This perspective helps to better understand the key methodologies and advancements and identify critical areas for future research. The eight key research directions are outlined as follows: